~~SOMEONE PLEASE HELP ME ON THIS ONE...I HAVE POSTED IT 3 TIMES ALREADY AND NOONE RESPONDED. GOD BLESS SOMEONE WHO HELPS~~
How much water can be held by a cylindrical tank with a radius of 12 feet and a height of 30 feet?
a) 13,564.8 cubic feet
b) 2,260.8 cubic feet
c) 33,912 cubic feet
d) 54,259.20 cuic feet

Answer by Earlsdon(6103)

To find the volume of a cylinder, you multiply the area of the base (a circle) by the height of the cylinder.
V = (pi)(r^2)(h) Substitute the given measurements of the cylinder.
V = (3.14)(12^2)(30)
V = (3.14)(144)(30)
V = 13,564.8 Cubic feet.